A manufacture wishes to design a hard disk with capacity of 60GB or more (using the standard definition of 1GB =230bytes).if the technology used to manufacture the disk allows 512bytes sectors, 1024sector/track,and 60%tracks/platter,how many platters are required
1
Expert's answer
2020-10-28T08:47:05-0400

As per the given question,

Capacity of hard disk "=60GB"

"1GB=2^{30}=1073741824" bytes

Number of bytes per sector "=512" bytes/sector

Number of sector per track "=1024" sector /track

Number of track per platter "=60\%" tracks /platter

The capacity of one platter "=512\times 1024\times \frac{60}{100}=314572.8"  bytes/platter

Hence,capacity of one platter in GB "=\frac{314572.8}{1073741824}=0.0002929GB"

Hence, number of required platter to make 60GB hard disk "=\frac{60}{0.0002929}=204848.0"